OpenStudy (anonymous):

Evaluate lim θ->0 (sin(3θ)cot(2θ))/(2θcot(θ))

OpenStudy (anonymous):

I think you first have to re-write it as: \[\lim_{\theta \rightarrow 0}\frac{\frac{1}{2\theta}cot(2\theta)}{\frac{1}{sin(3\theta)}cot(\theta)}\] Which is of the form \(\frac{\infty}{\infty}\) and use l'hopitals.

myininaya (myininaya):

\[\lim_{\theta \rightarrow 0}\frac{\sin(3 \theta)*\frac{\cos(2 \theta)}{\sin(2 \theta)}}{2 \theta*\frac{\cos( \theta)}{\sin(\theta)}}\] \[=\lim_{\theta \rightarrow 0}\frac{3 \theta \frac{\sin3 \theta}{3 \theta}*\cos(\theta)*\frac{2 \theta}{\sin(2 \theta)}}{2*\cos(\theta)*\frac{\theta}{\sin(\theta)}}\] \[=\frac{3*0*1*1*1}{2*1*1}\]

myininaya (myininaya):

oops i made a mistake

myininaya (myininaya):

do you see in the numerator that I multiply by 2 theta but i didn't divide by 2 theta

myininaya (myininaya):

so actually the limit is not 0

myininaya (myininaya):

so if we make that fix we should get 3/4 if i didnt make another mistake \[=\frac{3*1*1*1*\frac{1}{2}}{2*1*1}=\frac{\frac{3}{2}}{2}=\frac{3}{4}\]

myininaya (myininaya):

yep looks good you can also compare your answer graphically if you wish

OpenStudy (anonymous):

Hrm.. I'm assuming you used l'hopital's in that first step there, but I thought we had to have it in one of the specific forms. Where originally it's of the form \(\frac{0\times \infty}{0 \times \infty}\). I suppose you can split them up and take the derivative of each though rather than inverting the 0's as I did.

myininaya (myininaya):

no i didn't use l'hopital i just rewrote the function

myininaya (myininaya):

i wrote the cot as cos/sin

myininaya (myininaya):

and then i used the property that as x->0, sinx/x->1

myininaya (myininaya):

or i mean i put it in terms of sinx and cosx

myininaya (myininaya):

and cosx->1 as x->0

OpenStudy (anonymous):

oh I see. Bleh.

myininaya (myininaya):

\[\lim_{\theta \rightarrow 0}\frac{\sin(3 \theta)*\frac{\cos(2 \theta)}{\sin(2 \theta)}}{2 \theta \frac{\cos(\theta)}{\sin(\theta)}}\] \[=\lim_{\theta \rightarrow 0}\frac{3 \theta*\frac{\sin(3 \theta)}{3 \theta}*\cos(2 \theta)*\frac{2 \theta}{\sin(2 \theta)}*\frac{1}{2 \theta}}{2 \cos(\theta) \frac{\theta}{\sin(\theta)}}\]

myininaya (myininaya):

there maybe that looks better

myininaya (myininaya):

i went back in and put the thing i was missing lol

myininaya (myininaya):

perhaps lhopital way is better since this looks messey but i don't know lhopital could be messey too

OpenStudy (anonymous):

it probably is what with all those chain rules

OpenStudy (anonymous):

and product rules

myininaya (myininaya):

maybe this is the best way for this one

OpenStudy (anonymous):

probably.

OpenStudy (anonymous):

wow thank you guys :)